OpenStudy (anonymous):

The asymptotes of q(x) = 1/2 csc 2x are

OpenStudy (anonymous):

Rewrite this as: \[q(x)=\frac{1}{2 \sin(2x)} \] \[2 \sin(2x) = 0 \implies \sin(2x) = 0 \implies 2x=n \pi \implies x = \frac{n \pi}{2}; n \in \mathbb{Z}\]